Lab #13: Work and Angular Momentum Wear running sh
Answer #1

while climbing up a hill (smooth ) of some angle i

force = mg sini

Work = F*d = mgsini *d

Power = Work/time = mgsini*d/t

b) let us say that power = P W

to carry a mass work done = mg*sini*l

(so more is the mass, more is the work required to be done)

and the power is constant

so, 200kg requires takes more amount of time than 100kg

c) similarly 200kg requres more amount of work

d) Angular momentum = m*v*r

it is the magnitude of rotation of a body, which is the product of its moment of inertia and its angular velocity.